Solve.
5/2x + 1/4x = 7/4 + x.
The solution set is x =_____/

Answer by richwmiller(17219)

convert 5x/2 to 10x/4
and add like terms
11x/4=7/4+4x/4
subtract 4x/4 from both sides
7x/4=7/4
multiply both sides by 4
7x=7
x=1